OpenStudy (anonymous):

how do i find the axis of symmetry of y=x^2-4x=7?

OpenStudy (anonymous):

i mean y=x^2-4x+7

OpenStudy (anonymous):

This is a quadratic function, so the graph is a parabola. The axis of symmetry passes through the vertex. Do you know a way to find the vertex?

OpenStudy (anonymous):

i got (2,14) as the vertex

OpenStudy (anonymous):

its probably wrong though.. i did (b/2 , c(4/2) )

OpenStudy (anonymous):

well im guessing it would be 4/2 for x and then 7(4/2) for y

OpenStudy (anonymous):

Yeah, that doesn't seem to check out. Looks like you're using the short cut from the quadratic formula, the x-coordinate of the vertex is -b/2a

OpenStudy (anonymous):

that's what someone on here told me to do.

OpenStudy (anonymous):

so it would be -4/2(1)

OpenStudy (anonymous):

It's a quick and reliable method. Here, since you're looking for an equation of the line of symmetry, you don't need the y-coordinate of the vertex, just x=h : h = -b/2a

OpenStudy (anonymous):

h = -(-4)/(2(1)) = 2.

OpenStudy (anonymous):

i need to find the vertex too though

OpenStudy (anonymous):

Don't forget the negative sign. b = -4 not just 4.

OpenStudy (anonymous):

Well, then the y-coordinate of the vertex is the function evaluated at y=f(h).

OpenStudy (anonymous):

so the line of symmetry is 2? or that's just the x?

OpenStudy (anonymous):

i.e. take the x=2 you got for the axis of symmetry and put it into the equation to get y.

OpenStudy (anonymous):

The line of symmetry is x=2. You have to include the 'x=' because the line is defined by an equation.

OpenStudy (anonymous):

so y=7(2)?

OpenStudy (anonymous):

y=(2)^2-4(2)+7. You have to use the whole equation.

OpenStudy (anonymous):

i get it now!

OpenStudy (jdoe0001):

hmmm same as before

OpenStudy (jdoe0001):

did you find the vertex?

OpenStudy (anonymous):

im doing it right now

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

y=-1?

OpenStudy (anonymous):

Alternate method is to complete the square to put the function in vertex form.

OpenStudy (anonymous):

Check your signs . .

OpenStudy (jdoe0001):

notice the "a" component, that is, the leading coefficient is positive, that means is going upwards VERTICALLY, so the symmetry for the vertical one, will be the "x" coordinate of the vertex

OpenStudy (anonymous):

so it's not -1, it's positive?

OpenStudy (jdoe0001):

well \(\bf \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\\ -\cfrac{b}{2a} \implies -\cfrac{-4}{2(1)} \implies 2\)

OpenStudy (anonymous):

so it's (2,2)

OpenStudy (jdoe0001):

well, you don't need the y-coordinate, since the axis of symmetry is just the x-coordinate only

OpenStudy (anonymous):

ok. and then the vertex would be (2,1)? that's what i got

OpenStudy (jdoe0001):

\(\bf c-\cfrac{b^2}{4a} \implies 7-\cfrac{(-4)^2}{4(1)} \implies 7-\cfrac{\cancel{16}}{\cancel{4}} \implies 7-4\)

OpenStudy (anonymous):

so (2,3)?

OpenStudy (jdoe0001):

for the vertex, yes

OpenStudy (anonymous):

ok thanks again!

OpenStudy (jdoe0001):

yw